# Designing Consensus-Based Distributed Filtering over Directed Graphs

### v1

@@ 5 Simulations
Consider the target dynamics described by
$A=\left(\begin{array}[]{cc}I_{2}&TI_{2}\\ 0&I_{2}\end{array}\right),$ (48)
where $T=0.25$ is the discretization interval. The process noise covariance is defined as
@@
Two kinds of sensors are employed in the sensor network: position sensors and velocity sensors. The observation matrix for the position sensors is given by
$C_{p}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\end{array}\right),$ (50)
with the measurement noise covariance $R_{p}=\text{diag}\{1,1\}$. Similarly, the observation matrix for the velocity sensors is represented by
$C_{v}=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\end{array}\right),$ (51)
@@
with the measurement noise covariance $R_{v}=\text{diag}\{5,5\}$. In this sensor network, it is assumed that sensor 1, 2, and 4 are the position sensors, while sensor 3 and 5 serve as the velocity sensors. The initial state is set as $x_{0}=[1;1;1;1]$, and the initial state estimate is set as a random variable with the mean $\hat{x}_{i,k|k}=x_{0}$ and the initial estimation covariance $P_{i,0|0}=\text{diag}\{10,10,10,10\}$. The parameter $\mu_{ij}$ is selected as $\frac{1}{l_{ii}+a_{il}+1}$.
The mean square error (MSE) is utilized to evaluate the performance of the estimator based on the Monte Carlo method, described by
$MSE_{i,k}=\frac{1}{M}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,k}-x^{(l)}_{k}\|^{2}_{2},$ (52)
@@
where $M$ is the trial number, and $\hat{x}^{(l)}_{i,k}$ and $x^{(l)}_{k}$ are the state estimate and the true state at the $l$-th trial, respectively. In the simulation, $M$ is set as $1000$. All simulations are executed in MATLAB R2020a with an Intel Core i5-1135G7 CPU @ 2.40 GHz.
@@
Two examples are designed to verify the effectiveness of the proposed algorithm. Example 1 aims to demonstrate the steady-state performance of the proposed COMDF as the fusion step $l$ increases. Example 2 is designed to show the performances and properties of COMDF in comparison with other existing consensus-based distributed filers.
@@
Figure 2: Illustration figure for the steady-state performance of COMDF with the increasing fusion step $l$.
@@
Example 1: Fig. 2 exhibits the steady-state performance of COMDF of five sensors with the increasing fusion step $l$. It is shown that the performance gap between the centralized filter and COMDF is exponential convergence (Theorem 3). In addition, a small fusion step can also ensure the performance of COMDF.
@@
Example 2: To assess the performance of the proposed COMDF, three other algorithms are considered: the consensus-on-measurement distributed filter (CMDF) from [6], the consensus-on-information distributed filter (CIDF) from [4], and the centralized Kalman filter (CKF) in Section 4.1. For comparison, it is assumed that the communication topology is undirected in Fig. 1, and the fusion step is set as $l=10$.
Fig. 3 displays the performance of four algorithms with the increasing time step $k$. As the time step $k$ tends to infinity, four algorithms converge. CKF and COMDF are the algorithms that fix the gain matrix $K$, hence, they may not be optimal in every step. However, as $k$ tends to infinity, CKF approaches optimality. Meanwhile, the performance of COMDF approaches that of CKF with the increasing fusion step. Consequently, before the algorithms converge, CIDF and CMDF show a faster convergence velocity. CIDF adopts the covariance intersection method to handle the noise correlations, and the steady-state performance is degraded.
Figure 3: Illustration figure for the performance of four algorithms with the increasing time step $k$.
Table 1 presents the time consumption and the memory usage of four algorithms in Example 2. All four algorithms exhibit similar memory usage, but substantial differences emerge in terms of time consumption. Notably, COMDF demonstrates the smallest time consumption among the three distributed algorithms, attributed to its transmission of only measurements. Conversely, CIDF exhibits the highest time consumption among the three distributed filters, possibly due to a higher frequency of the inverse operation in the algorithm.
Table 1: The time consumption and the memory usage of four algorithms in Example 2.
Algorithm Time Consumption (s) Memory Usage (MB)
COMDF      23.864      2192
CMDF      32.709      2183
CIDF      144.438      2189
CKF      2.005      2194
6 Conclusions
@@
This paper proposes a consensus-on-measurement distributed filter over directed graphs, embedded with an augmented leader-following measurement fusion strategy. The parameters are designed to guarantee the uniformly upper bound of the estimation error covariances. The steady-state and transient performances are analyzed with the increasing fusion step, and the relations between the proposed distributed filter and the centralized filter are revealed. In the future, it is desired to design a dynamic gain matrix to optimize the distributed filter at every step, aiming for a faster convergence. Additionally, there is an intention to explore a new parameter design method to reduce the spectral radius of the communication matrix.
@@ Appendix A PROOF of Proposition 1
Then, the augmented vector $\varepsilon^{(l)}_{k}$ can be obtained as
$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (54)
where $G$ is given in (13). When $l=0$, it follows
$\varepsilon^{(0)}_{i,k}$ $=z^{(0)}_{i,k}-y_{k}$ (55)
$=CAe_{i,k-1}-C\omega_{k-1}-\nu_{k}.$
Similarly, the augmented vector $\varepsilon^{(0)}_{k}$ is
$\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})$ (56)
$~{}~{}~{}~{}-1_{N}\otimes\nu_{k}.$
Next, it can be concluded that
$\varepsilon^{(l)}_{k}$ $=G\varepsilon^{(l-1)}_{k}$ (57)
$=G^{l}\varepsilon^{(0)}_{k}.$
@@
Appendix B PROOF of Lemma 2
According to Lemma 2, since $\rho(G)<1$, it follows $\lim_{l\to+\infty}G^{l}=0$. Observing (17) and (18), $G^{l}$ exists in all terms. Hence, this lemma can be proven.
Appendix C PROOF of Proposition 2
Utilizing the state estimator equations (4) and (5) in conjunction with the dynamical system (1), the estimation error $e_{i,k}$ for sensor $i$ is
@@ Appendix C PROOF of Proposition 2
$e_{k}$ $=(I_{N}\otimes(A-KCA))e_{k-1}$ (58)
$~{}~{}~{}~{}+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)e_{k-1}$
@@
×$~{}~{}~{}~{}\times G^{l}(I_{N}\otimes C))(1_{N}\otimes\omega_{k-1})$
@@
By denoting the matrices as (24), (25), and (26), $e_{k}$ has the following form
$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}).$
Finally, $P_{k|k}$ can be calculated as (22) and (23) according to $P_{k|k}=E\{e_{k}e^{T}_{k}\}$
Appendix D PROOF of Lemma 6
Under the condition $0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}$ and the fact $l_{ij}\leq 0,i\neq j$, it holds
$1-\mu_{il}(l_{ii}+a_{il})\geq 0$ (59)
and
$-\mu_{il}l_{ij}\geq 0.$ (60)
By combining (27), (28), (59), and (60), it can be concluded that the matrix $G$ is nonnegative. By utilizing Lemma 3 and observing (29), it follows
$\rho(G)$ $\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}G_{ij}$
$\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}g_{[ij,l]}$
$\leq 1-\mu_{il}a_{il}.$
Hence, the conclusion is drawn that $\rho(G)\leq 1$.
Subsequently, it will be demonstrated that the eigenvalues of $G$ do not include the value $1$. By substituting the value $1$ into the characteristic polynomial of $G$, one has
$|I-G|=|\Lambda(\mathcal{L}\otimes I_{r}+B)|.$
@@ Appendix D PROOF of Lemma 6
Appendix E PROOF of Theorem 1
@@
Given Assumption 2, it is well known that the matrix $A-KCA$ is Schur stable [2]. Applying Lemma 2, as $l$ tends to infinity, $\lim_{l\to\infty}G^{l}\to 0$. Consequently, one observes $\lim_{l\to\infty}(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)\to 0$. It is no doubt that there exists a positive $l$ such that $\mathcal{A}(l)$ is Schur stable.
In matrix theory, it is established that the spectral radius of $\rho(\mathcal{A}(l))$ is bounded by the spectral norm $\|\mathcal{A}(l)\|_{2}$, i.e., $\rho(\mathcal{A}(l))\leq\|\mathcal{A}(l)\|_{2}$. Using the identity $\|J_{1}\otimes J_{2}\|_{2}=\|J_{1}\|_{2}\|J_{2}\|_{2}$, it follows
$\|\mathcal{A}(l)\|_{2}\leq\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}.$ (61)
Let $\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}<1$. After some algebraic manipulations, one has
$\|G\|^{l}_{2}<\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (62)
Since $10<\|G\|_{2}<$ 1, there exists
$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ (63)
such that when $l>l_{0}$, $\mathcal{A}(l)$ is Schur stable. Consequently, it can be proven that $P_{k|k}$ is uniformly upper-bounded. Since
@@
×$~{}~{}~{}~{}\times[0^{T}_{n\times n(i-1)},I^{T}_{n},0^{T}_{n\times n(N-i)}]^{T},$
@@
Appendix F PROOF of Theorem 2
Under Assumption 1, 2, and 3, it can be concluded that $\mathcal{A}(l)$ and $\mathcal{A}_{cc}$ are Schur stable. In addition, $\Phi(l)$ and $\Phi_{cc}$ are also uniformly upper-bounded due to the boundedness of $K$, $C$, $Q$, $R$, and $G^{l}$. By utilizing Theorem 1 in [7], it can be proven that $P_{k|k}$ and $P_{cc,k|k}$ converge to $P^{(l)}$ and $P_{cc}$, respectively.
Appendix G PROOF of Lemma 7
@@
$P^{(l)}-P_{cc}$ $=\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}$ (68)
$~{}~{}~{}~{}+\sum^{k-1}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
By performing an infinite number of iterations and utilizing the fact that $\mathcal{A}_{cc}$ is Schur stable, it holds
$\lim_{k\to\infty}\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}=0$
and
$P^{(l)}-P_{cc}=\sum^{\infty}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$
Next, similarly to (65), $\bar{\Phi}(l)$ in (67) can be calculated as (44).
Appendix H PROOF of Theorem 3
@@
By combining (69), (71), (74), and (75), a positive number $M_{1}$ can be designed such that
$\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$ (76)
@@
By combining (69), (74), and (77), a positive number $M_{2}$ can be obtained such that
$\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$ (78)
Appendix I PROOF of Lemma 8
Based on Assumption 3 and Lemma 2, as $l$ tends to infinity, one has $\lim_{k\to\infty}G^{l}=0$. By combining (24), (25), and (26) and comparing them to (38), (39), and (40), this lemma can be proven.
Appendix J PROOF of Theorem 4
Utilizing Lemma 8, (22), and (36), it follows
$\lim_{l\to+\infty}P_{k|k}=P_{cc,k|k}.$ (79)
Based on (64), this conclusion can be drawn.

### v2

@@ V Simulations
Consider the target dynamics described by $A=\text{diag}\{I_{2}~TI_{2};0~I_{2}\}$, where $T=0.25$ is the discretization interval. The process noise covariance is defined as
@@
Two kinds of sensors are employed in the sensor network: position sensors and velocity sensors. The observation matrix for the position sensors is given by $C_{p}=[I_{2}~0_{2}]$ with the measurement noise covariance $R_{p}=\text{diag}\{1,1\}$. Similarly, the observation matrix for the velocity sensors is represented by $C_{v}=[0_{2}~I_{2}]$ with the measurement noise covariance $R_{v}=\text{diag}\{5,5\}$. In this sensor network, sensors 1, 2, and 4 are designated as the position sensors, while sensors 3 and 5 serve as the velocity sensors. The initial state is set as $x_{0}=[1;1;1;1]$, and the initial state estimate is set as a random variable with the mean $\hat{x}_{i,k|k}=x_{0}$ and the initial estimation covariance $P_{i,0|0}=\text{diag}\{10,10,10,10\}$. The parameter $\mu_{ij}$ is selected as $\frac{1}{l_{ii}+a_{il}+1}$. The mean square error (MSE) is utilized to evaluate the performance of the estimator based on the Monte Carlo method, described by
$MSE_{i,k}=\frac{1}{M}\sum^{M}_{m=1}\|\hat{x}^{(m)}_{i,k}-x^{(m)}_{k}\|^{2}_{2},$ (32)
@@
where $M$ is the trial number, and $\hat{x}^{(m)}_{i,k}$ and $x^{(m)}_{k}$ are the state estimate and the true state at the $m$-th trial, respectively. In the simulation, $M$ is set as $1000$. All simulations are executed in MATLAB R2024b with an Intel Core Ultra 9 185H CPU @ 2.30 GHz.
@@
Two examples are designed to verify the effectiveness of the proposed algorithm. We refer to our proposed consensus-based distributed filter as CBDF. Example 1 aims to demonstrate the steady-state performance of the proposed filter as the fusion step number $l$ increases. Example 2 is designed to show the performances and properties in comparison with other existing distributed filers.
@@
Figure 3: Illustration figure for the steady-state performance of CBDF with the increasing fusion step number $l$.
@@
Example 1: Fig. 3 exhibits the steady-state performance of CBDF of five sensors with the increasing fusion step number $l$. It is shown that the performance gap between the centralized filter and CBDF is exponential convergence (Theorem 3). In addition, a small fusion step can also ensure the performance of CBDF.
Figure 4: The steady-state error of sensor 1 across 1000 Monte Carlo simulations for four algorithms. In each box plot, the red dotted line represents the mean, while the blue solid line indicates the median. The upper and lower edges of each box correspond to the 25th and 75th percentiles, respectively. Red circles are used to highlight outliers that fall beyond 1.5 times the interquartile range.
@@
Example 2: To assess the performance of the proposed CBDF, three other algorithms are considered: the consensus-on-information distributed filter (CIDF) from [5], asymptotically optimal distributed filter (CMDF) from [4], and the centralized Kalman filter (CKF) in Section IV-A. For comparison, it is assumed that the communication topology is undirected in Fig. 2, and the fusion step is set as $l=10$. The Metropolis weights are adopted in CIDF and CMDF. We define the steady-state error for sensor $i$ as
$E_{i,\infty}=\|\hat{x}_{i,\infty}-x_{\infty}\|^{2}_{2}.$ (33)
Fig. 4 shows the steady-state error of sensor 1 for four algorithms from 1000 Monte Carlo experiments. It is observed that CKF exhibits the best performance among the four algorithms, while CBDF demonstrates good performance. It is worth mentioning that this simulation is executed over undirected graphs for comparison, since the weight design methods are not well established for directed graphs [20].
Table I presents the execution time and memory usage of the four algorithms in Example 2. All four algorithms exhibit similar memory usage, but substantial differences emerge in terms of time consumption. The CBDF requires less computation time, owing to its reliance primarily on simple addition and subtraction operations.
TABLE I: The execution time and memory usage in Example 2.
Algorithm Execution Time (s) Memory Usage (MB)
CBDF 5.21 3294
CMDF 14.49 3341
CIDF 41.69 3295
CKF 1.07 3326
VI Conclusion
@@
This paper proposed a consensus-based distributed filter over directed graphs, embedded with an augmented leader-following information fusion strategy. The parameters were designed to guarantee the uniform upper bound on the estimation error covariances. The steady-state performance was analyzed as the fusion step number increases, and the relations between the proposed distributed filter and the centralized filter were revealed. In the future, we aim to design a dynamic gain matrix to optimize the distributed filter at each step, with the goal of achieving faster convergence. Additionally, we intend to explore a new parameter design method to reduce the spectral radius of the communication matrix.
@@ Appendix A PROOF of Proposition 1
Then, the augmented vector $\varepsilon^{(l)}_{k}$ can be obtained as $\varepsilon^{(l)}_{k}=G\varepsilon^{(l-1)}_{k}$. When $l=0$, it follows $\varepsilon^{(0)}_{i,k}=z^{(0)}_{i,k}-y_{k}=CAe_{i,k-1}-C\omega_{k-1}-\nu_{k}.$ Similarly, the augmented vector $\varepsilon^{(0)}_{k}$ is $\varepsilon^{(0)}_{k}=(I_{N}\otimes CA)e_{k-1}-(I_{N}\otimes C)(1_{N}\otimes\omega_{k-1})-1_{N}\otimes\nu_{k}.$ Next, it can be concluded that $\varepsilon^{(l)}_{k}=G\varepsilon^{(l-1)}_{k}=G^{l}\varepsilon^{(0)}_{k}$.
@@
Item 3): Utilizing the state estimator (6) in conjunction with the dynamical system (1), the estimation error $e_{i,k}$ for sensor $i$ is
@@
$e_{k}$ $=(I_{N}\otimes(A-KCA))e_{k-1}+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)e_{k-1}$ (34)
@@
$~~~~\times G^{l}(I_{N}\otimes C))(1_{N}\otimes\omega_{k-1})$
@@
By denoting the matrices as (8), (9), and (10), $e_{k}$ has the following form $e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}).$
Item 4): $P_{k|k}$ can be calculated as (11) and (12) according to $P_{k|k}=E\{e_{k}e^{T}_{k}\}$ .
Appendix B PROOF of Lemma 5
Consider the row sum of the block matrix, and it holds
$~~~\sum^{n}_{j=1}g_{[ij,l]}$ $=\Big(1-\mu_{il}a_{il}\Big)\otimes I_{r_{l}}.$ (35)
Under the condition $0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}$ and the fact $l_{ij}\leq 0,i\neq j$, it holds $1-\mu_{il}(l_{ii}+a_{il})\geq 0$ and $\mu_{il}l_{ij}\geq 0.$ By combining (13) and (14), it can be concluded that the matrix $G$ is nonnegative. By utilizing Lemma 3 and observing (35), it follows $\rho(G)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}G_{ij}\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}g_{[ij,l]}\leq 1-\mu_{il}a_{il}.$ Hence, the conclusion is drawn that $\rho(G)\leq 1$. Subsequently, it will be demonstrated that the eigenvalues of $G$ do not include the value $1$. By substituting the value $1$ into the characteristic polynomial of $G$, one has $|I-G|=|\Lambda(\mathcal{L}\otimes I_{r}+B)|.$
@@ Appendix B PROOF of Lemma 5
Appendix C PROOF of Theorem 1
By employing Lemma 2 and considering (8), it is evident that $\rho(\mathcal{A}(l))<1$ can guarantee the convergence of the distributed filter. The following results provide a lower bound of the fusion step number $l$.
@@ Appendix C PROOF of Theorem 1
Given Assumption 1, it is well known that the matrix AKCAA-KCA is Schur stable [2]. Applying Lemma 2, as $l$ tends to infinity, $\lim_{l\to\infty}G^{l}\to 0$. Consequently, one observes $\lim_{l\to\infty}(I_{N}\otimes K)G^{l}(I_{N}\otimes CA)\to 0$. It is no doubt that there exists a positive $l$ such that $\mathcal{A}(l)$ is Schur stable. In matrix theory, it is established that the spectral radius of $\rho(\mathcal{A}(l))$ is bounded by the spectral norm $\|\mathcal{A}(l)\|_{2}$, i.e., $\rho(\mathcal{A}(l))\leq\|\mathcal{A}(l)\|_{2}$. Using the identity $\|J_{1}\otimes J_{2}\|_{2}=\|J_{1}\|_{2}\|J_{2}\|_{2}$, it follows $\|\mathcal{A}(l)\|_{2}\leq\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}.$ Let $\|A-KCA\|_{2}+\|K\|_{2}\|G\|^{l}_{2}\|CA\|_{2}<1$. After some algebraic manipulations, one has
$\|G\|^{l}_{2}<\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$
Since $0<\|G\|_{2}<1$, there exists $l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ such that when $l>l_{0}$, $\mathcal{A}(l)$ is Schur stable. Consequently, it can be proven that $P_{k|k}$ is uniformly upper-bounded. Since
@@
$~~~~\times[0^{T}_{n\times n(i-1)},I^{T}_{n},0^{T}_{n\times n(N-i)}]^{T},$
@@
Appendix D PROOF of Theorem 2
Under Assumptions 1, 2, and 3, it can be concluded that $\mathcal{A}(l)$ and $\mathcal{A}_{cc}$ are Schur stable. In addition, $\Phi(l)$ and $\Phi_{cc}$ are also uniformly upper-bounded due to the boundedness of $K$, $C$, $Q$, $R$, and $G^{l}$. By utilizing Theorem 1 in [6], it can be proven that $P_{k|k}$ and $P_{cc,k|k}$ converge to $P^{(l)}$ and $P_{cc}$, respectively.
Appendix E PROOF of Lemma 6
@@
$P^{(l)}-P_{cc}$ $=\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}+\sum^{k-1}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$ (40)
By performing an infinite number of iterations and utilizing the fact that $\mathcal{A}_{cc}$ is Schur stable, it holds $\lim_{k\to\infty}\mathcal{A}^{k}_{cc}(P^{(l)}-P_{cc})(\mathcal{A}^{T}_{cc})^{k}=0$ and $P^{(l)}-P_{cc}=\sum^{\infty}_{j=0}\mathcal{A}^{j}_{cc}\bar{\Phi}(l)(\mathcal{A}^{T}_{cc})^{j}.$ Next, similarly to (37), $\bar{\Phi}(l)$ in (39) can be calculated as (28).
Appendix F PROOF of Theorem 3
@@
By combining (41), (43), (46), and (47), a positive number $M_{1}$ can be designed such that $\|P^{(l)}-P_{cc}\|_{2}\leq M_{1}l^{Nr}\rho(G)^{l-Nr}.$
@@
By combining (41), (46), and (48), a positive number $M_{2}$ can be obtained such that $\|P^{(l)}-P_{cc}\|_{2}\leq M_{2}\|G\|^{l}_{2}.$